# Spectrum radii of trees

### v1

@@ 3 Weighted Directed trees
Claim 1: For any positive integer $i$, $\frac{(i+1)\sqrt{4m+3}}{2i+1}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$ and $\frac{(2i+1)(m+1)}{i\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$.
@@
It remains to show that for any integer $i$ with $10\leq i\leq m-$ 1, $g(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$ implies that $g(i+1)\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$. Observe that
@@
Claim 3: For any integer $s$ with $32m\leq s\leq 4m+$ 3, $\frac{s}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$.
@@
Let $h(i)=\frac{2m+i}{\sqrt{4m+3}}$. By Claim 2, $h(0),h(1)\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$. For any integer $i$ with $0\leq i\leq m$, by Claim 1, $\frac{(m-i+1)\sqrt{4m+3}}{2(m-i)+1}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$, implying that
@@
Let $k$ be an integer with $k\geq 2$, and let $r_{0}=\left\lfloor{\frac{1}{4}k}\right\rfloor+2$. By Lemma 5.1 and Propositions 5.25.35.4 and 5.5, for each integer $s$ with $\left\lfloor{\frac{1}{4}k}\right\rfloor\leq s\leq k$, $\frac{s}{\sqrt{k}}\in{\mathscr{W}}_{r_{0}}(\sqrt{k})$ holds, implying that
@@
Then, by Corollary 4.5, for any integer $r$ in $[r_{0},k]$, there exists a tree $T$ with $\Delta(T)=r$ and $\lambda(T)=\sqrt{k}$. $\Box$
6 Problems and Remarks
@@
It can be proved by applying Proposition 4.4 that, for any positive integers $p$ and $q$, $\frac{\sqrt{p}+\sqrt{q}}{2}\in\Lambda$ if $4|(p-q)$, $\sqrt{p+\sqrt{q}}\in\Lambda$ if $p\geq q$, and $\frac{\sqrt{2(2p+1)+2\sqrt{4q+1}}}{2}\in\Lambda$ if $q\leq p^{2}+p$. As an example, we provide a proof for the conclusion that $\frac{\sqrt{p}+\sqrt{q}}{2}\in\Lambda$ whenever $4|(p-q)$.
@@
For any positive integers $p$ and $q$ with $p\geq q$, if $p-q$ is a multiple of $4$, then there exists a tree $T$ with $\Delta(T)=1+\max\{q,\frac{p-q}{4}\}$ and $\lambda(T)=\frac{\sqrt{p}+\sqrt{q}}{2}$.
@@
By Proposition 4.4, there exists a tree $T$ with $\Delta(T)=\max\{q,d\}+1$ and $\lambda(T)=\frac{\sqrt{p}+\sqrt{q}}{2}$. $\Box$
We wonder if $4|(p-q)$ is a necessary condition for $\frac{\sqrt{p}+\sqrt{q}}{2}$ to be a member of $\Lambda$. For any positive integers $s$ and $t$, $\frac{s+t}{2}=\frac{\sqrt{s^{2}}+\sqrt{t^{2}}}{2}$. As $P(G,\lambda)\in{\cal P}$, each rational root of $P(G,\lambda)=0$ is an integer. It follows that $\frac{s+t}{2}\in\Lambda$ if and only if $s$ and $t$ have the same parity, i.e., $4|(s^{2}-t^{2})$.
Problem 6.3.
Is it true that for any positive integers $p$ and $q$, $\frac{\sqrt{p}+\sqrt{q}}{2}\in\Lambda$ if and only if $p-q$ is a multiple of $4$?
Let $T$ be any tree with $\alpha=\lambda(T)$ and $r=\Delta(T)$. Then, (1) yields that $\frac{1}{4}\alpha^{2}+1<r\leq\alpha^{2}$. By definition, $\alpha\in{\mathscr{W}}_{r}(\alpha)$. By Theorem 1.1, $\alpha^{-1}\in{\mathscr{W}}_{r}(\alpha)$. We wonder if this property holds for all numbers in ${\mathscr{W}}_{r}(\alpha)$.
Problem 6.4.
Let $\lambda\in\Lambda$ and $r$ be an integer with $\frac{1}{4}\alpha^{2}+1<r\leq\alpha^{2}$. Is it true that $q^{-1}\in{\mathscr{W}}_{r}(\alpha)$ for every $q\in{\mathscr{W}}_{r}(\alpha)$?

### v2

@@ 1 Introduction
Claim 1: For any $i\in{\mathbb{N}}^{+}$, $\frac{(i+1)\sqrt{4m+3}}{2i+1}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$ and $\frac{(2i+1)(m+1)}{i\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$.
@@
It remains to show that for any $i\in\{0\}\cup\llbracket m-1\rrbracket$, $g(i)\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$ implies that $g(i+1)\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$. Observe that
@@
Claim 3: For any $s\in{\mathbb{N}}^{+}$ with $32m\leq s\leq 4m+$ 3, $\frac{s}{\sqrt{4m+3}}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$.
@@
Let $h(i)=\frac{2m+i}{\sqrt{4m+3}}$. By Claim 2, $h(0),h(1)\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$. For any $i\in\{0\}\cup\llbracket m\rrbracket$, by Claim 1, $\frac{(m-i+1)\sqrt{4m+3}}{2(m-i)+1}\in{\mathscr{W}}_{m+2}(\sqrt{4m+3})$, implying that
@@
Let $k$ be an integer with $k\geq 2$, and let $r_{0}=\left\lfloor{\frac{1}{4}k}\right\rfloor+2$. By Lemma 4.1 and Propositions 4.24.34.4 and 4.5, for each integer $s$ with $r_{0}\leq s\leq k$, $\frac{s}{\sqrt{k}}\in{\mathscr{W}}_{r_{0}}(\sqrt{k})\subseteq{\mathscr{W}}_{s}(\sqrt{k})$ holds. Note that
@@
By Corollary 3.3, for any integer $s$ in $[r_{0},k]$, there exists a tree $T$ with $\Delta(T)=s$ and $\lambda(T)=\sqrt{k}$. $\Box$
5 Further study
@@
It can be proved by applying Theorem 1.1 that, for any $p,q\in{\mathbb{N}}^{+}$, $\frac{\sqrt{p}+\sqrt{q}}{2}\in\Lambda$ if $4|(p-q)$, $\sqrt{p+\sqrt{q}}\in\Lambda$ if $p\geq q$, and $\frac{\sqrt{(2p+1)+\sqrt{4q+1}}}{\sqrt{2}}\in\Lambda$ if $q\leq p^{2}+p$. As an example, we provide a proof for the conclusion that $\frac{\sqrt{p}+\sqrt{q}}{2}\in\Lambda$ whenever $4|(p-q)$.
@@
For any $p,q\in{\mathbb{N}}^{+}$ with $p\geq q$, if $p-q$ is a multiple of $4$, then there exists a tree $T$ with $\Delta(T)=1+\max\left\{q,\frac{p-q}{4}\right\}$ and $\lambda(T)=\frac{\sqrt{p}+\sqrt{q}}{2}$.
@@
By Corollary 3.4, there exists a tree $T$ with $\Delta(T)=\max\{q,d\}+1$ and $\lambda(T)=\frac{\sqrt{p}+\sqrt{q}}{2}$. $\Box$
We wonder if $4|(p-q)$ is a necessary condition for $\frac{\sqrt{p}+\sqrt{q}}{2}$ to be a member of $\Lambda$.
Problem 5.3.
Is it true that for any $p,q\in{\mathbb{N}}^{+}$, $\frac{\sqrt{p}+\sqrt{q}}{2}\in\Lambda$ if and only if $4\big{|}(p-q)$?
Let $T$ be any tree with $\alpha=\lambda(T)$ and $r=\Delta(T)$. Then, (1) yields that $\frac{1}{4}\alpha^{2}+1<r\leq\alpha^{2}$. By definition, $\alpha\in{\mathscr{W}}_{r}(\alpha)$. From the proving of Lemma 2.3, $\alpha^{-1}\in{\mathscr{W}}_{r}(\alpha)$. We wonder if this property holds for all numbers in ${\mathscr{W}}_{r}(\alpha)$.
Problem 5.4.
Let $\lambda\in\Lambda$ and $r\in{\mathbb{N}}^{+}$ with $\frac{1}{4}\alpha^{2}+1<r\leq\alpha^{2}$. Is it true that $q^{-1}\in{\mathscr{W}}_{r}(\alpha)$ for every $q\in{\mathscr{W}}_{r}(\alpha)$?
Note that some similar work on Laplacian spectral radii of trees has been completed in [8].
Conflict of interest: The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability statements: Not applicable.